OpenStudy (anonymous):

An athlete swings a ball, connected to the end of a chain, in a horizontal circle. The athlete is able to rotate the ball at the rate of 8.54 rev/s when the length of the chain is 0.600 m. When he increases the length to 0.900 m, he is able to rotate the ball only 6.31 rev/s. What is the centripetal acceleration of the ball at 8.54 rev/s?

OpenStudy (jamesj):

What's the formula for centripetal acceleration for uniform circular motion? If v = speed of ball (meters/second) r = radius of motion (meters) w (omega) = angular velocity (radians/second) then the acceleration, a, is given by \[ a = \frac{v^2}{r} = \omega^2 r \] So either of the two expressions can be used. Remember also that \( v = \omega r \).

OpenStudy (anonymous):

we know basically, that arch lenghth=angle subtend by this at centre * radius of the circle. L=\[\theta \times radius\]=>\[v(velocity)=\omega \times radius\] after differentiate on both side w.r.to time. and centripetal acceleration of the body is \[a(centripetal acceleration)=v(linear velocity)^2\div radius OR \omega(angular velocity)^2\times radius\].apply the formula you get your answer.......